OpenStudy (anonymous):

Find the ratio of the following geometric sequence. 4, -2, 1,-1/2 ,1/4 , ...

OpenStudy (cwrw238):

-2 / 4 or if you like (-1/2 ) / 1

OpenStudy (anonymous):

Just choose one value and divide it by the value that comes before it, so, Ratio = -2/4 = -1/2 or 1/-2 = -1/2 Or -1/2 / 1 = -1/2

OpenStudy (anonymous):

i got -1/8

OpenStudy (cwrw238):

no -2 / 4 = -1/2 is the correct ratio